At a bakery there are 45 different kinds of breads and sweets. If the ratio of sweets to breads is 4:5, how many types of breads are there?
Answer by TimothyLamb(4379)

x = breads
y = sweets
---
x + y = 45
y/x = 4/5
---
x + y = 45
y = (4/5)x
---
put the system of linear equations into standard form
---
x + y = 45
(4/5)x - y = 0
---
x + y = 45
0.8x - y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = breads = 25
y = sweets = 20
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---